OpenStudy (anonymous):

What are the foci of the graph x^2-4y^2=16

OpenStudy (anonymous):

\[ \frac{x^2}{4^2}-\frac{y^2}{2^2}=1 \]

OpenStudy (precal):

yes

OpenStudy (anonymous):

a=4 b=2 c^2= a^2 + b^2 = 20

OpenStudy (anonymous):

\[ c= 2 \sqrt{5} \]

OpenStudy (anonymous):

Elias, my son. You are a very thorough educator. I have learnt that you have been educating for 40 years.

OpenStudy (anonymous):

The foci are (c,0) and (-c,0)

OpenStudy (anonymous):

I cannot be as good as Sir Isaac Newton !!!!